OpenStudy (anonymous):

812/1000 in simplest form

OpenStudy (anonymous):

what is 812/1000 in simplest form?!i'm only in 6th grade,i don't know all this stuff

OpenStudy (anonymous):

\[\frac{ 812 }{1000 }\]

OpenStudy (anonymous):

HELP!!!!!!!!!!!!

OpenStudy (anonymous):

You can just simplify by halving every time until you can't simplify anymore \[\frac{ 406 }{ 500 }\] \[\frac{ 203 }{ 250 }\]